OpenStudy (anonymous):

a wooden block is dropped from the top of a building 100m high above ground simultaneously a bullet is fired from the foot of the building upwards witha velocity of 100m/s the height above the ground at which they meet is?

OpenStudy (anonymous):

can anyone solve this

OpenStudy (anonymous):

the ans=95.1m

OpenStudy (anonymous):

Remember the kinematic equation? The bullet has position 0, initial velocity 100 m/s and acceleration -10m/s^2. The block has 0 initial velocity, initial position -100 and acceleration 10m/s^2. So: 100t - (1/2)(10)t^2 = 100 - (1/2)(10)t^2. So: t = 1 Plug in on the bullet equation: S = 100(1) - 5(1) = 95

OpenStudy (ajprincess):

let the height travelled by the wooden block be s1 s1=1/2*10*t^2 let the height travelled by the bullet be s2 s2=100t-1/2*10*t^2 both the cases time same s1+s2=100 100=5t^2+100t-5t^2 100=100t t=1 s1=5*1 =5m. s2=100-5 =95m

OpenStudy (anonymous):

There is a minor rounding due to the fact that I took g as 10 m/s^2 instead of 9.81 or something like that.